### v3

# Sentiment Reasoning for Healthcare

Abstract
Transparency in AI healthcare decision-making is crucial for building trust among AI and users. Incorporating reasoning capabilities enables Large Language Models (LLMs) to understand emotions in context, handle nuanced language, and infer unstated sentiments. In this work, we introduce a new task - Sentiment Reasoning - for both speech and text modalities, along with our proposed multimodal multitask framework and dataset. Sentiment Reasoning is an auxiliary task in sentiment analysis where the model predicts both the sentiment label and generates the rationale behind it based on the input transcript. Our study conducted on both human transcripts and Automatic Speech Recognition (ASR) transcripts shows that Sentiment Reasoning helps improve model transparency by providing rationale for model prediction with quality semantically comparable to humans while also improving model performance (1% increase in both accuracy and macro-F1) via rationale-augmented fine-tuning. Also, no significant difference in the semantic quality of generated rationales between human and ASR transcripts. All code, data (English-translated and Vietnamese) and models are published online: https://github.com/leduckhai/MultiMed.
[…]
Sentiment Reasoning for Healthcare
Khai-Nguyen Nguyen∗1,8, Khai Le-Duc∗2,3,8, Bach Phan Tat4, Duy Le5, Long Vo-Dang6, Truong-Son Hy7,8 1College of William and Mary, USA 2University of Toronto, Canada 3University Health Network, Canada 4KU Leuven, Belgium 5Bucknell University, USA 6University of Cincinnati, USA 7 University of Alabama at Birmingham, USA 8 FPT Software AI Center, Vietnam knguyen07@wm.edu, duckhai.le@mail.utoronto.ca, thy@uab.edu
**footnotetext: Equal contribution
1 Introduction
Figure 1: Visualized pipeline for Sentiment Reasoning. Given an input transcript (either human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label). It comprises of two tasks: Sentiment Classification and Rationale Generation.
In recent years, speech sentiment analysis has emerged as a significant interdisciplinary field at the intersection of natural language processing (NLP), machine learning, and automatic speech recognition (ASR). This field focuses on the automated detection and interpretation of human emotions and attitudes conveyed through speech, overcoming the limitations of text-based sentiment analysis Murugaiyan and Uyyala (2023) by analyzing the attitude expressed through human voice. The global market for sentiment analysis is projected to expand from an estimated value of US\$4 billion in 2023 to US\$10.1 billion by 2030, exhibiting a compound annual growth rate (CAGR) of 14.2% over the forecast period from 2023 to 2030 Inc (2024).
Sentiment analysis plays a pivotal role across diverse domains, with healthcare being particularly significant. In healthcare customer service, it facilitates real-time evaluation of customer satisfaction, enhancing empathetic and responsive interactions Xia et al. (2009); Na et al. (2012). Moreover, sentiment analysis aids in monitoring the emotional well-being of patients Cambria et al. (2012a), including those with mental health issues such as suicide Pestian et al. (2012). However, these studies only work on text-based sentiment analysis rather than speech-based sentiment analysis.
Despite its potential, speech sentiment analysis presents several technical challenges. First, speech signals are inherently noisy Rajnoha and Pollák (2011) and exhibit significant variability due to differences in accents Turan et al. (2020), speaking styles Shafran and Rose (2003), and recording conditions Yamagishi et al. (2010), complicating the extraction of reliable acoustic features Chen et al. (2017). Second, emotions are subjective Wearne et al. (2019), complex Golan et al. (2006), and multidimensional, making accurate categorization difficult even for humans Kuusikko et al. (2009), thereby necessitating the role of explainable artificial intelligence (AI). Third, given the critical nature of healthcare decisions, where errors can have severe consequences, transparency in AI decision-making is essential to build trust among machines, healthcare professionals, and patients Antoniadi et al. (2021).
To tackle these challenges, we introduce a novel multimodal framework for a novel task: Sentiment Reasoning. Reasoning in AI is crucial for sentiment analysis because it enables deeper understanding beyond surface-level sentiment polarity. By incorporating reasoning capabilities, AI can contextualize emotions within broader discourse, account for nuanced expressions, handle ambiguous language, and infer underlying intent or sentiment that may not be explicitly stated. This enhances the accuracy and reliability of sentiment analysis by considering complex human communication factors that simplistic methods may overlook. Our contributions are as follows:
  1. 1.
    We introduce a new task: Sentiment Reasoning for both speech and text modalities, along with our dataset MultiMed-SA
  2. 2.
    We propose our novel multimodal speech-text Sentiment Reasoning framework
  3. 3.
    We empirically evaluate the baselines on our dataset using state-of-the-art backbone models
  4. 4.
    We provide in-depth analysis of rationale / Chain-of-Thought (CoT)-augmented training
All code, data (English-translated and Vietnamese) and models are published online [^1] https://github.com/leduckhai/MultiMed.
2 Data
2.1 Data Collection
The dataset employed for constructing the Sentiment Reasoning dataset was VietMed Le-Duc (2024), the largest and publicly accessible medical ASR dataset in the world. We then annotated sentiment labels (POSITIVE, NEUTRAL, or NEGATIVE) and their corresponding rationales (the reason for this label). This dataset comprises real-world doctor-patient conversations covering a broad spectrum of all available ICD-10 codes (all medical topics), wherein doctors diagnose conditions, explain medical issues, and provide treatment advice.
Details of data annotation pipeline, annotation guidelines, data quality control, and data statistics are shown Section B in the Appendix.
2.2 Data Statistics
Split Label Count Percentage
Neutral 2844 49.94%
Train Negative 1694 29.74%
Positive 1157 20.32%
Neutral 958 43.88%
Test Negative 701 32.11%
Positive 524 20.01%
Table 1: Distribution of sentiment labels in the dataset.
Table 1 shows the distribution of sentiment labels in the dataset. In total, the dataset comprises 7878 samples with sentiment labels with rationales. This distribution reflects the dataset’s slight emphasis on neutral content, typical in medical conversations involving detailed explanations and advice.
3 Sentiment Reasoning Framework
3.1 Informal Definition
As shown in Figure 1, in Sentiment Reasoning, given an input transcript (either human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label). It comprises of two tasks: Sentiment Classification and Rationale Generation.
Previous works Wadhwa et al. (2024); Chen et al. (2024); Hsieh et al. (2023); Ho et al. (2022) have shown that rationale-augmented targets consistently improve the performance of Language Models (LMs). Furthermore, having the option to output the model’s rationale during inference is highly desirable in many scenarios.
3.2 Formal Definition
Let $x^{T}_{1}:=x_{1},x_{2},...,x_{T}$ be an audio signal of length $T$. Let $C$ be the set of all possible sentiment classes, we should build a speech Sentiment Reasoning model $f$ that both estimates the probability $p(c|x^{T}_{1})$ for each $c\in C$ and generates its rationale sequence $r^{M}_{1}$ of $M$ length.
The decision rule to predict a single sentiment class is:
$x^{T}_{1}\to\hat{c}=\arg\max_{c\in C}f(c|x^{T}_{1})$ (1)
The decision rule to generates the corresponding rationale sequence is:
$x^{T}_{1}\to r^{M}_{1}=\arg\max_{r^{*}}h(r^{*}|x^{T}_{1})$ (2)
For text-based Sentiment Reasoning, the input audio signal $x^{T}_{1}$ could be replaced with a word sequence (human transcript) $w^{N}_{1}$ of length N, thus ASR model is not needed.
3.3 ASR Model
An ASR model aims to convert audio signal into text by mapping an audio signal $x^{T}_{1}$ to the most likely word sequence $w^{N}_{1}$. The relation $w^{*}$ between the acoustic and word sequence is:
$w^{*}=\operatorname{arg}\max_{w_{1}^{N}}\,p(w_{1}^{N}|x_{1}^{T})$ (3)
By utilizing Bayes’ Theorem, the probability $p(x)$ can be ignored during maximization since it merely serves as a normalization factor and does not affect the outcome:
$\begin{split}p(w_{1}^{N}|x_{1}^{T})&=\frac{p(x_{1}^{T}|w_{1}^{N})p(w_{1}^{N})}{p(x_{1}^{T})}\\ &\propto p(x_{1}^{T}|w_{1}^{N})p(w_{1}^{N})\end{split}$ (4)
Therefore:
$w^{*}=\operatorname{arg}\max_{w_{1}^{N}}\underbrace{p(x_{1}^{T}|w_{1}^{N})}_{\text{acoustic model}}\cdot\underbrace{p(w_{1}^{N})}_{\text{language model}}$ (5)
3.4 Language Model for Sentiment Reasoning
3.4.1 Sentiment Classification
Let the transcribed audio signal (ASR transcript) $w^{N}_{1}$ serve as the input for the Sentiment Classification model $g$, which maps $w^{N}_{1}$ to a class label $\hat{c}$:
$w^{N}_{1}\to\hat{c}=\arg\max_{c\in C}g(c|w^{N}_{1})$ (6)
$g$ is trained to minimize a loss function $\mathscr{L}(g(w^{N}_{1}),\hat{c})$. The optimal parameters $\theta$ of the model are found by solving the optimization problem $\min_{\theta}\mathscr{L}(g(w^{N}_{1};\theta),\hat{c})$. Once trained, the model can predict the class of the transcribed audio signal by evaluating $\hat{c}=g(w^{N}_{1})$.
3.4.2 Rationale Generation
Let the transcribed audio signal (ASR transcript) $w^{N}_{1}$ serve as the input for the Rationale Generation model $h$, which maps $w^{N}_{1}$ to a rationale sequence $r^{M}_{1}$ of $M$ length:
$w^{N}_{1}\to r^{M}_{1}=\arg\max_{r^{*}}h(r^{*}|w^{N}_{1})$ (7)
$h$ is trained to minimize a loss function $\mathscr{L}(h(w^{N}_{1}),r^{M}_{1})$. The optimal parameters $\theta$ of the model are found by solving the optimization problem $\min_{\theta}\mathscr{L}(g(w^{N}_{1};\theta),r^{M}_{1})$. Once trained, the model can generate rationale of the transcribed audio signal by evaluating $r^{M}_{1}=h(w^{N}_{1})$.
4 Experimental Setups
4.1 ASR Model
We employed hybrid ASR setup using wav2vec 2.0 encoder Le-Duc (2024) to transcribe speech to text. First, we generated alignments obtained by using Gaussian-Mixture-Model/Hidden-Markov-Model (GMM/HMM) as labels for wav2vec 2.0 Baevski et al. (2020) neural network training. The labels used in the acoustic modeling are context-dependent phonemes, triphones in this case. In GMM/HMM process, we used a CART (Classification And Regression Tree) Breiman (2017) to tie the states $s$, resulting 4501 CART labels:
$\begin{split}&p(x_{1}^{T}|w_{1}^{N})=\sum_{[s_{1}^{T}]}\prod_{t=1}^{T}p(x_{t},s_{t}|s_{t-1},w_{1}^{N})\\ &=\sum_{[s_{1}^{T}]}\prod_{t=1}^{T}\underbrace{p(s_{t}|s_{t-1},w_{1}^{N})}_{\text{transition prob.}}\cdot\underbrace{p(x_{t}|s_{t},s_{t-1},w_{1}^{N})}_{\text{emission prob.}}\end{split}$ (8)
After inputting CART labels for hybrid wav2vec 2.0 training, we employed frame-wise cross-entropy (fCE) loss Good (1952) to train the acoustic model.
To transcribe speech given the acoustic observations, the acoustic model and n-gram language model Ney et al. (1994) should be combined based on the Bayes decision rule using Viterbi algorithm Forney (1973) which recursively computes the maximum path to a find best-path in the alignment graph of all possible predicted words to the acoustic observations:
$\begin{split}w_{1}^{N}&=\operatorname{arg}\max_{N,w_{1}^{N}}p\Bigl{(}\prod_{n=1}^{N}p(w_{n}|w_{n-m}^{n-1})\\ &\cdot\max_{[s_{1}^{T}]}\prod_{t=1}^{T}p(x_{t},s_{t}|s_{t-1},w_{1}^{N})\Bigr{)}\end{split}$ (9)
Finally, acoustic model and n-gram language model pruning (beam search) is used to only focus on the most promising predicted words at each time step $t$ Ortmanns et al. (1997).
The final ASR model has 118M trainable parameters and Word-Error-Rate (WER) of 29.6% on VietMed test set.
4.2 Language Model for Sentiment Reasoning
4.2.1 Encoder
The encoder architecture is naturally well-suited for Sentiment Classification, which can be reformulated into the classical classification task. To this end, we directly apply a linear classifier to the output of the encoders. However, encoders can not generate rationales.
We use phoBERT (110M params) Nguyen and Nguyen (2020), a version of RoBERTa Liu et al. (2019) pre-trained on 20GB Vietnamese text, and ViHealthBERT (110M params) Minh et al. (2022), phoBERT trained on 32GB of Vietnamese text in the healthcare domain. For ViHealthBERT, we report the syllable version which achieved better performance than the word version.
4.2.2 Generative Models
[…]
Encoder-Decoder: BARTpho (139M params) Tran et al. (2022a) is the Vietnamese variant of BART Lewis et al. (2019) trained on 20GB of Vietnamese text from Wikipedia and news corpus. ViT5 (223M params) Phan et al. (2022) is the Vietnamese version of T5 Raffel et al. (2020) trained on 71GB of Vietnamese text from CC100 Conneau et al. (2019).
Decoder: We use Vistral-7B-Chat Nguyen et al. (2023) and vmlu-llm [^2] https://huggingface.co/vtrungnhan9/vmlu-llm. Both models have Mistral-7BJiang et al. (2023a) as their backbone. These models were chosen based on their performance on the vmlu benchmark (Vietnamese Multitask Language Understanding) [^3] https://vmlu.ai/leaderboard which compares the performance of Vietnamese Large Language Models (LLMs) on various language understanding tasks.
4.3 Training with Rationale
Our rationale-augmented training methods are based on, to our knowledge, the current state-of-the-art CoT-distillation approaches for each architecture.
(i) Multitask Training Hsieh et al. (2023): We train our encoder-decoders using distilling step-by-step. Distilling step-by-step is a multitask training approach that prepends particular prefixes to the input, guiding the model to output either the answer or generate a rationale. Hsieh et al. found that it consistently improves encoder-decoders performance compared with single-task training which treats rationale and label predictions as a single task.
(ii) Post-thinking Chen et al. (2024): For decoder-based models, we augment the training targets by append the human rationale to the label (¡LABEL¿ ¡RATIONALE¿) in a single prompt. Previous works have shown that post-thinking achieved impressive performance Chen et al. (2024); Wadhwa et al. (2024) and compared to pre-thinking where the model first generates its chain-of-thought then provide the label (¡RATIONALE¿ ¡LABEL¿), post-thinking is more stable and token-efficient Chen et al. (2024); Wadhwa et al. (2024) as the model suffers less from hallucination, consistently yields better performance and is more resource efficient as users can already retrieve the target label from the first generated token.
4.4 Rationale Format
We further study the effects of the format of the rationale on the performance of the generative models. In particular, given the human rationale and human label, we further prompt GPT-3.5-turbo [^4] https://platform.openai.com/docs/models to enhance the rationale into two different format:
  • Elaborated rationale: An elaborated version of the human rationale that is 1-2 sentence(s) long, grounded on the provided human rationale and the sentiment label.
  • CoT rationale: A step-by-step, elaborated version of the human rationale, which includes the following steps: (1) identifies the medical entity, (2) extracts the progress of the corresponding medical entity in the transcript, and (3) provides the elaborated rationale on the sentiment grounded on the provided human rationale, the sentiment label, and information from steps (1) and (2). This approach is inspired by aspect-based sentiment instruction-tuning approaches Varia et al. (2022).
4.5 Evaluation Metrics
For Sentiment Classification task, we employ both accuracy and F1 score, the harmonic mean of precision and recall. It is particularly useful for evaluating models, as it considers both false positives and false negatives, providing a more comprehensive performance metric than accuracy alone.
For Rationale Generation, we employ ROUGE (Recall-Oriented Understudy for Gisting Evaluation) score Lin (2004), which is a set of metrics used to evaluate the quality of generated text by comparing the overlap of n-grams Ney et al. (1994), word sequences, and word pairs between the generated texts and reference texts. Also, we employ BERTScore Zhang et al. , which is an evaluation metric that leverages pre-trained BERT embeddings Devlin et al. (2019) to compute cosine similarity scores between generated texts and reference texts, capturing contextual and semantic nuances more effectively than traditional count-based metrics like ROUGE. BERTscore correlates well with human judgment on sentence-level as proved by Zhang et al. .
5 Results and Analysis
Model Acc. F1 Neg. F1 Neu. F1 Pos. Mac F1 R-1 R-2 R-L R-Lsum BERTscore
Encoder (Label Only)
PhoBERT 0.6674 0.6969 0.6607 0.6377 0.6651
ViHealthBERT 0.6752 0.6970 0.6718 0.6535 0.6741
Encoder-Decoder (Label Only)
ViT5 0.6628 0.6922 0.6687 0.6007 0.6545
BARTpho 0.6523 0.6870 0.6571 0.5841 0.6427
Decoder (Label Only)
vmlu-llm 0.6592 0.6768 0.6769 0.5911 0.6483
Vistral7B 0.6716 0.6858 0.6771 0.6398 0.6676
Encoder-Decoder (Label + Rationale)
ViT5 0.6633 0.6936 0.6572 0.6335 0.6615 0.3910 0.2668 0.3653 0.3660 0.8093
BARTpho 0.6619 0.7029 0.6460 0.6265 0.6585 0.3871 0.2613 0.3658 0.3683 0.8077
Decoder (Label + Rationale)
vmlu-llm 0.6729 0.7039 0.6714 0.6307 0.6687 0.3947 0.2467 0.3789 0.3796 0.8086
Vistral7B 0.6812 0.7152 0.6765 0.6425 0.6781 0.4155 0.2788 0.3880 0.3900 0.8101
Table 2: Baseline performance of encoders, encoder-decoders, and decoders on the Vietnamese human transcript. From left to right is: Accuracy, F1-{negative, neutral, positive, macro}, ROUGE-{1, 2, L, Lsum}, BERTscore. The Label Only models are models trained only with the label, serving as the baseline, while Label + Rationale indicates models trained with rationale. As the Label Only models are not trained to generate rationale, we do not evaluate them on ROUGE and BERTscore.
Model Acc. F1 Neg. F1 Neu. F1 Pos. Mac F1 R-1 R-2 R-L R-LSum BERTscore
Encoder (Label Only)
PhoBERT 0.6166 0.6418 0.6231 0.5658 0.6102
ViHealthBERT 0.6198 0.6307 0.6261 0.5934 0.6167
Encoder-Decoder (Label Only)
ViT5 0.6157 0.6412 0.6258 0.5523 0.6064
BARTpho 0.6056 0.6364 0.6156 0.5311 0.5944
Decoder (Label Only)
vmlu-llm 0.6216 0.6296 0.6551 0.5186 0.6011
Vistral7B 0.6299 0.6377 0.6537 0.5609 0.6174
Encoder-Decoder (Label + Rationale)
ViT5 0.6189 0.6305 0.6286 0.5837 0.6143 0.3571 0.2202 0.3350 0.3366 0.8044
BARTpho 0.6129 0.6523 0.6028 0.5665 0.6072 0.3956 0.2652 0.3728 0.3774 0.8106
Decoder (Label + Rationale)
vmlu-llm 0.6395 0.6585 0.6557 0.5723 0.6289 0.3853 0.2386 0.3663 0.3671 0.8092
Vistral7B 0.6354 0.6485 0.6479 0.5892 0.6285 0.3558 0.2237 0.3343 0.3394 0.7994
Table 3: Baseline performance of encoders, encoder-decoders, and decoders on the Vietnamese ASR transcript. Further information about our metrics can be found in Table 2.
Model Acc. F1 Neg. F1 Neu. F1 Pos. Mac F1
Encoder-Decoder (Label + Rationale)
ViT5_human 0.6633 0.6936 0.6572 0.6335 0.6615
ViT5_elaborate 0.6661 0.6903 0.6799 0.5985 0.6562
ViT5_cot 0.6619 0.6968 0.6552 0.6237 0.6586
BARTpho_human 0.6619 0.7029 0.6460 0.6265 0.6585
BARTpho_elaborate 0.6564 0.7031 0.6528 0.5870 0.6476
BARTpho_cot 0.6464 0.6922 0.6611 0.5287 0.6273
Decoder (Label + Rationale)
Vistral7B_human 0.6812 0.7152 0.6765 0.6425 0.6781
Vistral7B_elaborate 0.6688 0.6846 0.6647 0.6564 0.6685
Vistral7B_cot 0.6706 0.6725 0.6807 0.6477 0.6670
vmlu-llm_human 0.6729 0.7039 0.6714 0.6307 0.6687
vmlu-llm_elaborate 0.6867 0.7203 0.6868 0.6353 0.6808
vmlu-llm_cot 0.6821 0.6966 0.6779 0.6711 0.6819
Table 4: Performance of generative models on the different rationale formats on our test set. Human/elaborate/CoT specifies the format of rationale the model was trained on. Details in Section 4.4
Figure 2: Confusion matrix of the predicted classes versus the actual labels on human transcript, obtained from Vistral7B trained with human rationale
We evaluate and analyze our models performance on Table 2. Based on the obtained results, we make the following observations:
1. Encoders are efficient yet effective sentiment classification baselines: Encoder models yields the best performance compared to their encoder-decoder and decoder counterparts, with high accuracy scores (¿ 0.665) and stable F1 scores (macro F1 of both models ¿ 0.665). Furthermore, compared to other models, encoders are more parameter efficient which makes them highly economical for training and deployment. We further observe that domain-specific encoders yield notably better performance, with ViHealthBERT outperforming phoBERT in accuracy (+0.8%) and macro F1 (+0.9%).
2. Rationale-augmented training improve model performance: Consistent with previous findings, performing CoT-augmented training on both encoder-decoders and decoders improve our models performance compared to the baseline. Aside from ViT5, other models seem to have benefited from rationale-augmented training, especially the decoder-based LLMs. We observe an average accuracy gain of approximately +0.8% and average macro-F1 gain of approximately +1.3% for these models compared to their label-only counterparts. We further conducted a Student’s t-test Student (1908) and found that the gains are statistically significant for $\alpha=0.1$. This pattern holds for the results in Table 4. We observe a decline in all of our models performance on ASR data which is anticipated due to its WER of 29.6 %. Nonetheless, the models trained with rationale perform noticeably better than models without, with an average accuracy gain of +0.85% and macro F1 gain of +1.4%.
3. The format of post-thinking rationale doesn’t affect the generative models performance: We study the effects of the format of post-thinking rationale on the performance of generative models on Table 4 and observe that it is unclear whether there is a performance gain from more elaborated rationales. This result agrees with previous findings Wadhwa et al. (2024). Particularly, we observe that some models trained on the elaborated rationales have better performance (ViT5 and vmlu-llm). However, for the other models, there is a notable decline in all metrics.
4. Models are likely to misclassify POSITIVE and NEGATIVE transcripts as NEUTRAL: We study the confusion matrix of our best model on human transcript, Vistral7B finetuned with human rationale, on Figure 2. We observe a notable misclassification tendency between NEUTRAL and the other two classes (23.43% and 27.08% with NEGATIVE and POSITIVE respectively). On the other hand, we found that models can easily distinguish NEGATIVE transcripts from POSITIVE ones. This reflects the ambiguity of sentiment analysis data, as in certain cases there is not enough distinctive features that differentiate neutral from other sentiments. Furthermore, given the slightly imbalanced nature of our dataset with fewer POSITIVE examples, its average F1 score is the lowest among the three labels across all models.
5. Analysis of Generated Rationale: Compared to human rationale, we observe from Table 2 and Table 3 that the models trained with rationale have high BERTscore (around 0.8) with low ROUGE score, indicating that while the vocabulary used in the rationale is different, the overall semantic of the generated rationale remains similar to that of humans. Also, no noticeable changes in the semantic quality of rationale between human transcripts and ASR transcripts because BERTScore is still about 0.8 on both settings.
6 Conclusion
In this work, we introduce a new task - Sentiment Reasoning - for both speech and text modalities, along with the framework and dataset. In Sentiment Reasoning, given an input transcript (either human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label). It comprises of two tasks: Sentiment Classification and Rationale Generation.
We thoroughly evaluate the use of rationale during training to improve our models’ interpretability and performance. We found that rationale-augmented training improves model performance in Sentiment Classification in both human and ASR transcripts (1% increase in both accuracy and macro-F1). We also found that the generated rationales generally have different vocabulary to that of human but with similar semantics. Finally, we found no significant difference in the semantic quality of generated rationales between human and ASR transcripts.
7 Limitations
This study utilized the hybrid ASR system, which is generally recognized as superior in performance compared to the attention-based encoder-decoder or end-to-end ASR systems Lüscher et al. (2019); Prabhavalkar et al. (2023); Raissi et al. (2023). However, the hybrid ASR requires multiple steps, beginning with acoustic feature extraction and progressing through GMM-HMM modeling before transitioning to DNN-HMM modeling, which complicates reproducibility for non-experts.
Additionally, we employed a cascaded speech Sentiment Reasoning approach. This approach uses a previously trained ASR model to generate transcripts that are subsequently input into a LM for downstream sentiment classification and reasoning tasks. Consequently, the weights in the ASR model remain unchanged while the LM weights are updated. In spoken language processing, cascaded approach is prefered due to its straightforwardness, simplicity and superior accuracy Lu (2023); Bentivogli et al. (2021); Tran et al. (2022b); Tseng et al. (2023). Future work should consider the end-to-end sentiment analysis task, where weights in both the ASR model and LM are updated simultaneously, as it holds promise for improved performance.
8 Acknowledgement
We would like to thank Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for giving insightful feedback.
Appendix A Related Works
A.1 Multimodal Speech Sentiment Analysis
It is widely known that there have been two research directions in the field of speech sentiment analysis, as also confirmed by Chen et al. (2020).
  • Single modality model (unimodal): In speech sentiment analysis, single modality models focus on utilizing a single type of data to predict sentiment. These models may rely exclusively on acoustic features, such as pitch, tone, and rhythm, to infer emotional states from spoken language Li et al. (2019, 2018); Wu et al. (2019); Xie et al. (2019). Alternatively, they might use raw waveforms Tzirakis et al. (2018); Zheng et al. (2022); Villatoro-Tello et al. (2021) or the textual content of transcripts to predict sentiment Lakomkin et al. (2019). The strength of single modality models lies in their simplicity and specialization, allowing them to hone in on specific attributes of the data source they are designed for. However, this specialization can also be a limitation, as these models might miss out on the richer, more nuanced information that can be gleaned from combining multiple data types. Despite this, single modality models remain a fundamental approach in the field, providing valuable insights and serving as a benchmark for more complex multimodal systems.
  • Multimodality models: In speech sentiment analysis, multimodality models leverage the combined strengths of both acoustic and textual data to provide more accurate and nuanced sentiment predictions. While traditional models might rely solely on either the acoustic features—such as tone, pitch, and rhythm—or the text derived from speech transcripts, multimodal models integrate these two data streams. This integration allows for a more holistic understanding of sentiment, as it captures the emotional cues present in the speaker’s voice along with the contextual and semantic content of the spoken words. By maximizing the mutual information between these modalities, multimodal models can better discern subtleties in speech that single modality models might miss, leading to accuracy improvements Kim and Shin (2019); Cho et al. (2018); Gu et al. (2018); Eskimez et al. (2018); Zhang et al. (2019).
Our dataset is ideal for both single modal and multimodal research, as it includes both acoustic and text features.
A.2 ASR-based Speech Sentiment Analysis
Speech sentiment analysis on ASR transcripts is a field that aims to interpret and classify sentiments conveyed in spoken language. As technology advances, ASR systems have become increasingly proficient at transcribing spoken words into text with high accuracy Schneider et al. (2019); Baevski et al. (2020, 2019); Wang et al. (2021b); Chen et al. (2022); Wang et al. (2021a), providing a rich source of data for sentiment analysis. Sentiment analysis algorithms then analyze the transcribed text from speech signal, utilizing language models as decoders to detect positive, negative, or neutral sentiments Lu et al. (2020); Shon et al. (2021); Wu et al. (2022); Tashev and Emmanouilidou (2019); Kaushik et al. (2017).
In the era of deep learning, as surveyed by Al-Qablan et al. (2023), many researchers have been applying deep learning methods to the sentiment analysis process on transcript, leading to the development of various models like Convolutional Neural Networks (CNN), Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM), and Bidirectional LSTM (BLSTM) Araque et al. (2017); Devipriya et al. (2020); Yadav and Vishwakarma (2020). CNNs, primarily used for image processing, have been adapted for text by treating sentences as sequences of words and applying convolutional filters to capture local features. This approach helps in identifying crucial patterns within the text that are indicative of sentiment Kumar and Malarvizhi (2020); Wang et al. (2020). On the other hand, RNNs are designed to handle sequential data by maintaining a hidden state that captures the history of previous inputs, making them suitable for understanding the context and temporal dependencies in sentences. However, traditional RNNs face challenges with long-term dependencies due to issues like vanishing gradients, which is where LSTMs come in. LSTMs, an advanced form of RNNs, address these issues by incorporating gates that regulate the flow of information, allowing them to maintain and update long-term dependencies effectively. Furthermore, BLSTMs enhance this by processing the input sequence in both forward and backward directions, thus capturing dependencies from both past and future contexts simultaneously. This bidirectional approach is especially useful for sentiment analysis, where the interpretation of a word can depend heavily on both preceding and succeeding words. Together, these architectures provide powerful tools for sentiment analysis, each contributing unique strengths that can be leveraged depending on the specific requirements and characteristics of the data at hand Gandhi et al. (2021); Pal et al. (2018); Srinivas et al. (2021).
Developed by Google, BERT (Bidirectional Encoder Representations from Transformers) Kenton and Toutanova (2019) revolutionized NLP tasks by enabling models to understand the context of words in a sentence more effectively through its bidirectional training approach. Unlike previous models that read text input sequentially, BERT reads the entire sequence of words at once, capturing the full context and nuances of language. This capability allows BERT to excel in sentiment analysis, where understanding the subtleties of human emotion and opinion is paramount Alaparthi and Mishra (2020); Deepa (2021). BERT’s pre-training on vast amounts of text data, followed by fine-tuning on specific sentiment analysis tasks, further enhances its performance. By leveraging its powerful language representations, BERT can handle the complexities of sentiment analysis, such as sarcasm, idiomatic expressions, and context-dependent sentiment shifts, making it a preferred choice for applications ranging from social media monitoring to customer feedback analysis. The model’s ability to generalize across various domains and languages also contributes to its widespread adoption, offering robust and scalable solutions for sentiment analysis in diverse settings Hoang et al. (2019); Xu et al. (2019); Sousa et al. (2019); Alaparthi and Mishra (2021).
A.3 Speech Sentiment Analysis in Healthcare
Sentiment analysis in healthcare is an emerging field that leverages NLP and machine learning techniques to analyze and interpret the emotional tone conveyed in biomedical textual data. This technology is particularly useful for understanding patient feedback, monitoring public health trends, and improving patient-provider communication. By analyzing large volumes of data from sources such as social media, online reviews, electronic health records (EHRs), and patient surveys, sentiment analysis can provide valuable insights into patient experiences, satisfaction levels, and overall public sentiment towards healthcare services and policies. For instance, analyzing patient reviews on healthcare platforms can help identify common concerns and areas needing improvement, allowing healthcare providers to address issues proactively and enhance the quality of care. Additionally, sentiment analysis can play a critical role in mental health monitoring by detecting signs of distress or dissatisfaction in patient communications, enabling timely intervention and support. As this technology continues to evolve, it holds the promise of transforming healthcare by fostering a more patient-centric approach, enhancing service delivery, and ultimately improving patient outcomes Denecke and Deng (2015). However, the sentiments expressed in clinical narratives have not been extensively analyzed or exploited, based on the total number of previous works we have identified to the best of our knowledge:
  • Sentiment analysis from the medical web: Most sentiment analysis research in the medical domain focuses on web data, such as medical blogs and forums, to mine patient opinions or assess quality Ali et al. (2013); Xia et al. (2009); Na et al. (2012); Sokolova et al. (2013); Biyani et al. (2013); Ofek et al. (2013); Smith and Lee (2012); Sharif et al. (2014); Melzi et al. (2014).
  • Sentiment analysis from biomedical literature: In addition to the analysis of medical social media data, biomedical literature has been examined concerning the outcomes of medical treatments. Within this framework, sentiment denotes the results or efficacy of a treatment or intervention Niu et al. (2005); Sarker et al. (2011).
  • Sentiment analysis from medical text (except biomedical literature): Several researchers have focused on leveraging supplementary sources of medical texts to implement sentiment analysis and emotion detection methodologies, suicide notes or patient questionnaire for example Pestian et al. (2012); Cambria et al. (2012a); Liu and Singh (2004); Cambria et al. (2012b).
To the best of our knowledge, no literature among those cited has addressed speech sentiment analysis specifically within the domain of healthcare.
Appendix B Details about Data
B.1 Data Annotation
We use LLM pre-labeling as it helps speed up the labeling process through providing the annotators with the initial sentiment labels and the corresponding rationales. In the relabeling process, annotators go through each sample and inspect it manually. If the annotators deem the label and the rationale is appropriate, they can quickly move to the next sample. If not, the annotators can update the label and rationale to be more appropriate.
The data annotation process is as followed. First, all the subtitles are separated into different chunks. These segments are subsequently input into gpt-3.5-turbo, which conducts a weakly supervised 3-label classification task to categorize each segment as NEGATIVE, NEUTRAL, or POSITIVE. In addition to the sentiment label, gpt-3.5-turbo also provides a brief synthetic rationales for the classification, such as ’Negative medical condition’ or ’Objective description’. The labels and rationales generated by gpt-3.5-turbo are subsequently reviewed and independently corrected by a team of 3 developers.
B.2 LLM Prompt for Pre-labeling
gpt-3.5-turbo Annotate the sentiment (neutral, positive or negative) of the following sentence and provide a very short justification. The procedure is as follows: 1. If the segment shows clear emotional signs, annotate based on these signs. 2. If no emotional markings are present, determine if the segment is an objective description. Positive for beneficial facts/features, negative for detrimental facts/features, and neutral otherwise. 3. If not objective, check if there’s a preference expression. Positive for likes or positive views, negative for dislikes or negative views, and neutral if no preference is expressed. 4. If too short to determine sentiment, label as neutral. {3 in-context learning examples}
B.3 Annotation Guidelines
The definition of ”sentiment” encompasses both ”emotions” and ”facts” in our work. Existing works Chen et al. (2020) use both emotions and facts for sentiment labeling. To ensure consistency, our TESOL-certificated professional linguist has developed an initial guideline inspired by Chen et al. (2020) and revised it frequently if necessary as followed:
The NEGATIVE label is for chunks that discuss negative, serious diseases, disorders, symptoms, risks, negative emotions, or counter-positive statements (e.g. ”This would NOT bring a good outcome”). It also applies to incomplete chunks where the amount of negativity is greater than the amount of positivity.
The NEUTRAL label is for incomplete chunks where the ratio of negativity is equal to the ratio of positivity, as well as chunks that describe processes, ask questions, provide advice, or are too short.
The POSITIVE label is for chunks that discuss positive outcomes, recovery processes, positive emotions, or counter-negative statements (e.g. ”This will reduce discrimination”). It also applies to incomplete chunks where the ratio of positivity is greater than the ratio of negativity.
It is important to note that all chunks are considered independent, even though they may be incomplete and related to preceding or following chunks. Given that this data is derived from spoken language, the chunks contain a significant amount of filler words, which are disregarded in the labeling process. The majority of the NEUTRAL labels are attributed to chunks that involve sharing advice or descriptions. Additionally, the presence of modal verbs (e.g., should, would, need) often indicates advice sharing, thereby classifying the chunk as NEUTRAL regardless of its content.
B.4 Data Quality Control
During the independent annotation process conducted by three annotators, we observed a low inter-annotator agreement, a common occurrence in real-world datasets as noted by Chen et al. (2020). To address this issue, we implemented an alternative label merging approach. We convened a discussion meeting involving the three annotators and two reviewers (one professional linguist and one with a biomedical background). Each annotator was required to justify their chosen sentiment label and its corresponding rationale. A label and its rationale were selected based on the consensus of all three annotators and two reviewers, rather than a majority vote, as employed in other studies Aziz and Dimililer (2020); Saleena et al. (2018).
B.5 Data Statistics Discussion
The NEUTRAL category is the most predominant, accounting for a significant portion of the dataset. With 3802 instances for both train and test set, NEUTRAL sentiments make up approximately half of the dataset. This prevalence of NEUTRAL sentiment is expected, as also seen by a real-world conversational dataset Chen et al. (2020), given the nature of medical consultations, which often involve objective descriptions, explanations, and advice. The NEGATIVE category is the second most common, with around 2395 instances. NEGATIVE sentiments include discussions about serious diseases, negative emotions, and adverse medical outcomes. The substantial presence of negative sentiments reflects the medical context, where discussions about illnesses and symptoms are common. The POSITIVE category, while the least common, still represents a significant portion of the dataset with 1681 instances. POSITIVE sentiments typically involve discussions about recovery processes, positive outcomes, and favorable emotions.
B.6 Data Samples
Table 5 shows 9 examples with 3 samples per sentiment label in our dataset.
Transcript ENG Translation Label Rationale
bệnh nhân sẽ có những cái rối loạn
về mặt cảm xúc đôi khi có những
bệnh nhân đã rơi vào trạng thái trầm
cảm và đôi khi
The patient will suffer from emotional
disorder and sometimes depression
NEG. Emotional disorder
não đột quỵ đó thì nó liên quan đến
việc hình thành các cục máu đông
và việc cục máu đông đã nó trôi ra
là đi
Stroke is related to the formation of
blood clots and the fact that these
blood clots travel
NEG.
Negative medical
condition
nhầm lẫn với một cái nhóm thuốc
khác đó là nhóm thuốc gọi là thuốc
chống tiểu cầu tiểu cầu mà cụ
It’s often confused with
antiplatelet drugs
NEG. Confusion
điểm cần thiết phải lưu tâm rõ ràng
là cái người là bị béo phì đó
A crucial point is that the
overweight patient
NEU. Sharing advice
ra đó là cái hormone cortisol trong
máu cũng như là hormone về
catecholamine nó
The cortisol hormone in blood as well
as catecholamine
NEU.
Objective description
of hormones
có thể gọi đây là thuốc lẫn máu
hay là một số cái tên khác mà thì
nó có thể
You could call these blood-thinning
drugs or other names, and it can
NEU. Objective description
của nó không có cao nhưng mà rất
là hình thức thì rất là may mắn là
những năm gần đây thì mình có
một cái nhóm thuốc khác
It is not expensive, luckily, in recent
years there are another group of
medicine
POS. Expressing luck
để mà giảm xóa bỏ cái chuyện
hình thành cái cục máu đông đó
hiện ta sẽ dùng một số biện pháp
trong đó thì chủ
To reduce and eliminate the formation
of these blood clots, we use several
measures, one of which is
POS.
Avoid forming
blood clots
nhóm thuốc này á thì nó là rất là
lâu đời và nó không có mất tiền
rất là rẻ là
This group of drugs has been around
for a very long time and is very
cheap, with no cost
POS.
Long-standing and
inexpensive medication
Table 5: 9 examples with 3 samples per sentiment label and its corresponding rationale
Appendix C Details about Experimental Setups
C.1 Training Setup
Our encoders and encoder-decoders were trained on a cluster of 2 NVIDIA A40s with 46 GBs of memory. All models were trained on 30 epochs with with a learning rate of $-52e\text{-}$ 5 and batch size of 64. We evaluated every epoch with early stopping with patience = 3.
For the decoder-based LLMs, due to their massive number of parameters, we use LoRA Hu et al. (2021) for fine-tuning with hyperparameters: the rank of the update matrices $r=8$, and the LoRA scaling factor $\alpha=3$. We train our LLMs for 5 epochs with learning rate 2e-4.
We use the best model checkpoints for evaluation. Note that we do not perform hyperparameter tuning as we only aim to provide the initial benchmark results as well as studying the effects of CoT-augmented finetuning.
C.2 Student’s T-Test
A Student’s t-test, is a statistical method used to compare the means of one or two populations through hypothesis testing. It can assess whether a single group mean differs from a known value (one-sample t-test), compare the means of two independent groups (independent two-sample t-test), or determine if there is a significant difference between paired measurements (paired or dependent samples t-test). Figure 3 below is the code for reproducing Student’s t-test experiments.
Figure 3: Python code for reproducing Student’s t-test experiments
Appendix D Results on English subset
We use Google Translate to translate our data into English. To ensure that our translated data is usable, we randomly sampled 50 transcripts and check their quality. We further train English models on this English subset of our dataset to ensure full usability.
The result of our experiments is in Table 6. More information on the models used can be found in the same table. Overall, we found that rationale-augmented training also help boost the model’s performance. This finding is consistent with what when observed in our experiments in Section 5.
Model Acc. F1 Neg. F1 Neut. F1 Pos. Mac F1
Encoder (Label Only)
mBERT Devlin et al. (2018) 0.6001 0.5972 0.6320 0.5408 0.5900
BERT Devlin et al. (2018) 0.6143 0.6338 0.6245 0.5653 0.6079
Encoder-Decoder (Label Only)
mT0 Muennighoff et al. (2022) 0.6216 0.6303 0.6418 0.5670 0.6130
Flan-T5 Chung et al. (2022) 0.6157 0.6295 0.6385 0.5462 0.6048
Encoder-Decoder (Label + Rationale)
mT0 Muennighoff et al. (2022) 0.6175 0.6495 0.6253 0.5535 0.6094
Flan-T5 Chung et al. (2022) 0.6326 0.6487 0.6390 0.5978 0.6285
LLM (Label only)
Mistral7B Jiang et al. (2023b) 0.6290 0.6536 0.6322 0.5850 0.6236
LLM (Label + Rationale)
Mistral7B Jiang et al. (2023b) 0.6454 0.6768 0.6364 0.6176 0.6436
Table 6: Baseline performance of encoders, encoder-decoders, LLMs on the English human transcript. Further information about our metrics can be found in Table 2.
Appendix E Error Analysis
Transcript Label Pred. Human Rationale Model Rationale
VI: trả lại cho họ chất lượng cuộc
sống bình thường như bao
người khác là được nghe được
nói thế nhưng điều kỳ diệu đã
ENG: give them back a normal
quality of life like everyone
else, but a miracle has
happened
NEU. POS.
Mô tả khách quan
(Objective
description)
chất lượng cuộc sống
bình thường
(normal quality of life)
VI: những chia sẻ vô cùng hữu
ích và thiết thực vừa rồi ạ
có thể thấy là hầu hết người
bệnh nằm điều trị trong
ENG: with the extremely useful
and practical shares shared
just now, it can be seen that
most of the patients are in
hospital for treatment)
NEU. POS.
Mô tả khách quan
(Objective
description
thông tin hữu ích và
thiết thực
(useful and practical
information)
VI: khám suốt tiểu đường nó
vẫn mệt mỏi vô khám tai
biến bộ não vô khám nhưng
mà xương thì nó loãng
xương rất là nhiều
ENG: even after being examined
for diabetes, she still feels
tired, has had a stroke, and
has not been examined for
stroke, but her bones have
a lot of osteoporosis
NEU. NEG.
Mối quan tâm và
vấn đề sức khỏe
(Health concerns
and problems)
triệu chứng tiêu cực
của bệnh tiểu đường
và loãng xương
(negative
symptoms of
diabetes and
osteoporosis)
Table 7: Some misclassified transcripts from our best model with high confidence (¿0.99). VI means the Vietnamese transcript, EN means the transcript translated to English
We report our best model’s misclassified transcripts with the highest label confidence (defined as the softmax of the logits of the model prediction) in Table 7. By analyzing at the model’s rationale, we hypothesize that the model is confounded by the appearance of certain keywords that elicit either extremely positive ( hữu ích (helpful)) or negative, disease-related words and sentiment which pushes the model away from the NEUTRAL label.

### v4

@@
 Abstract
-Transparency in AI healthcare decision-making is crucial for building trust among AI and users.
-Incorporating reasoning capabilities enables Large Language Models (LLMs) to understand emotions in context, handle nuanced language, and infer unstated sentiments.
-In this work, we introduce a new task - Sentiment Reasoning - for both speech and text modalities, along with our proposed multimodal multitask framework and dataset.
+Transparency in AI healthcare decision-making is crucial.
+By incorporating rationales to explain reason for each predicted label, users could understand Large Language Models (LLMs)’s reasoning to make better decision.
+In this work, we introduce a new task - Sentiment Reasoning - for both speech and text modalities, and our proposed multimodal multitask framework and the world’s largest multimodal sentiment analysis dataset.
 Sentiment Reasoning is an auxiliary task in sentiment analysis where the model predicts both the sentiment label and generates the rationale behind it based on the input transcript.
-Our study conducted on both human transcripts and Automatic Speech Recognition (ASR) transcripts shows that Sentiment Reasoning helps improve model transparency by providing rationale for model prediction with quality semantically comparable to humans while also improving model performance (1% increase in both accuracy and macro-F1) via rationale-augmented fine-tuning.
+Our study conducted on both human transcripts and Automatic Speech Recognition (ASR) transcripts shows that Sentiment Reasoning  helps improve model transparency by providing rationale for model prediction with quality semantically comparable to humans while also improving model’s classification performance (+2% increase in both accuracy and macro-F1) via rationale-augmented fine-tuning.
 Also, no significant difference in the semantic quality of generated rationales between human and ASR transcripts.
-All code, data (English-translated and Vietnamese) and models are published online: https://github.com/leduckhai/MultiMed.
-[…]
-[…]
+All code, data (five languages - Vietnamese, English, Chinese, German, and French) and models are published online.
 Sentiment Reasoning for Healthcare
-Khai-Nguyen Nguyen∗1,8, Khai Le-Duc∗2,3,8, Bach Phan Tat4, Duy Le5, Long Vo-Dang6, Truong-Son Hy7,8 1College of William and Mary, USA 2University of Toronto, Canada 3University Health Network, Canada 4KU Leuven, Belgium 5Bucknell University, USA 6University of Cincinnati, USA 7 University of Alabama at Birmingham, USA 8 FPT Software AI Center, Vietnam knguyen07@wm.edu, duckhai.le@mail.utoronto.ca, thy@uab.edu
-**footnotetext: Equal contribution
+Khai-Nguyen Nguyen∗1, Khai Le-Duc∗2,3, Bach Phan Tat4, Duy Le5, Long Vo-Dang6, Truong-Son Hy7 1College of William and Mary, USA 2University of Toronto, Canada 3University Health Network, Canada 4KU Leuven, Belgium 5Bucknell University, USA 6University of Cincinnati, USA 7University of Alabama at Birmingham, USA [Uncaptioned image]knguyen07@wm.edu       [Uncaptioned image]duckhai.le@mail.utoronto.ca [Uncaptioned image]https://github.com/leduckhai/Sentiment-Reasoning
+(*)(*)footnotetext: Equal contribution
 1 Introduction
@@
 Given an input transcript (either human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label).
-It comprises of two tasks: Sentiment Classification and Rationale Generation.
-In recent years, speech sentiment analysis has emerged as a significant interdisciplinary field at the intersection of natural language processing (NLP), machine learning, and automatic speech recognition (ASR).
-This field focuses on the automated detection and interpretation of human emotions and attitudes conveyed through speech, overcoming the limitations of text-based sentiment analysis Murugaiyan and Uyyala (2023) by analyzing the attitude expressed through human voice.
-The global market for sentiment analysis is projected to expand from an estimated value of US\$4 billion in 2023 to US\$10.1 billion by 2030, exhibiting a compound annual growth rate (CAGR) of 14.2% over the forecast period from 2023 to 2030 Inc (2024).
-Sentiment analysis plays a pivotal role across diverse domains, with healthcare being particularly significant.
+It comprises of two tasks: (1) Sentiment Classification and (2) Rationale Generation .
+Traditional sentiment analysis only includes Sentiment Classification task, while our framework generates corresponding rationale to explain the reason behind each predicted sentiment label.
+9 examples with sentiment labels and their corresponding rationales in our dataset are shown in Table 7 in the Appendix.
+Sentiment analysis plays a pivotal role within the healthcare domain.
 In healthcare customer service, it facilitates real-time evaluation of customer satisfaction, enhancing empathetic and responsive interactions Xia et al.
@@
 (2012).
-However, these studies only work on text-based sentiment analysis rather than speech-based sentiment analysis.
+However, these studies only work on text-only sentiment analysis instead of speech-based sentiment analysis.
 Despite its potential, speech sentiment analysis presents several technical challenges.
-First, speech signals are inherently noisy Rajnoha and Pollák (2011) and exhibit significant variability due to differences in accents Turan et al.
-(2020), speaking styles Shafran and Rose (2003), and recording conditions Yamagishi et al.
-(2010), complicating the extraction of reliable acoustic features Chen et al.
+First, emotions conveyed through speech are subjective Wearne et al.
+(2019), complex Golan et al.
+(2006), and dependent on speaking styles Shafran and Rose (2003), making accurate sentiment classification difficult even for humans Kuusikko et al.
+(2009), thereby necessitating the role of explainable artificial intelligence (AI).
+Second, given the critical nature of healthcare decisions, where errors can have severe consequences, transparency in AI decision-making is essential to build trust among machines, healthcare professionals, and patients Antoniadi et al.
 (2021).
-Second, emotions are subjective Wearne et al.
-(2019), complex Golan et al.
-(2006), and multidimensional, making accurate categorization difficult even for humans Kuusikko et al.
-(2009), thereby necessitating the role of explainable artificial intelligence (AI).
-Third, given the critical nature of healthcare decisions, where errors can have severe consequences, transparency in AI decision-making is essential to build trust among machines, healthcare professionals, and patients Antoniadi et al.
-(2021).
-To tackle these challenges, we introduce a novel multimodal framework for a novel task: Sentiment Reasoning.
-Reasoning in AI is crucial for sentiment analysis because it enables deeper understanding beyond surface-level sentiment polarity.
-By incorporating reasoning capabilities, AI can contextualize emotions within broader discourse, account for nuanced expressions, handle ambiguous language, and infer underlying intent or sentiment that may not be explicitly stated.
-This enhances the accuracy and reliability of sentiment analysis by considering complex human communication factors that simplistic methods may overlook.
+To tackle challenges above, reasoning in AI is crucial for sentiment analysis because it enables deeper understanding beyond surface-level sentiment polarity via the textual explanations.
+Recent works on Chain-of-Thought (CoT) distillation Wadhwa et al.
+(2024); Chen et al.
+(2024); Hsieh et al.
+(2023); Ho et al.
+(2022) have revealed that training generative small language models (SLMs) on rationale-augmented targets (the CoT from larger models is provided along side with the target label) can help the SLM (1) perform better and (2) acquire the ability to generate rationale.
+Our work leverage these findings and prepare a set of human-labeled rationale to train our sentiment analysis models to do Rationale Generation and enhance their performance (Section 4.4 and 4.5).
+By incorporating rationales to explain reason for each predicted sentiment label, users could understand the model’s reasoning, facilitating better decision-making based on the classification results.
+Therefore, we introduce a novel multimodal framework for a novel task: Sentiment Reasoning, which comprises of two tasks: (i) Sentiment Classification, in which the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE), and (ii) Rationale Generation, in which the model generates rationale (the free-form text that explains reason for this label).
 Our contributions are as follows:
   1. 1.
-    We introduce a new task: Sentiment Reasoning for both speech and text modalities, along with our dataset MultiMed-SA
+    We introduce a new task: Sentiment Reasoning  for both speech and text modalities, along with the world’s largest multimodal sentiment analysis dataset, supporting five languages (Vietnamese, English, Chinese, German, and French)
   2. 2.
@@
     We provide in-depth analysis of rationale / Chain-of-Thought (CoT)-augmented training
-All code, data (English-translated and Vietnamese) and models are published online [^1] https://github.com/leduckhai/MultiMed.
+All code, data and models are published online.
 2 Data
 2.1 Data Collection
-The dataset employed for constructing the Sentiment Reasoning dataset was VietMed Le-Duc (2024), the largest and publicly accessible medical ASR dataset in the world.
+The dataset employed for constructing the Sentiment Reasoning  dataset was VietMed Le-Duc (2024), a large and publicly accessible medical ASR dataset.
+The dataset comprises real-world doctor-patient conversations.
 We then annotated sentiment labels (POSITIVE, NEUTRAL, or NEGATIVE) and their corresponding rationales (the reason for this label).
-This dataset comprises real-world doctor-patient conversations covering a broad spectrum of all available ICD-10 codes (all medical topics), wherein doctors diagnose conditions, explain medical issues, and provide treatment advice.
-Details of data annotation pipeline, annotation guidelines, data quality control, and data statistics are shown Section B in the Appendix.
+We then manually translate the transcripts from Vietnamese into other four languages: English, Chinese (Simplified and Traditional), German, and French, making the dataset six times larger.
+The full dataset (with 5 languages) includes 30000 samples, making it the largest multimodal sentiment analysis dataset, to the best of our knowledge (see Table 2).
+Our paper focuses mainly on the Vietnamese subset (Section 5) and the English subset (Appendix D).
+2.2 Data Annotation
+The annotation task consists of two primary steps.
+First, annotators are required to perform Sentiment Classification.
+Second, annotators are instructed to provide a rationale behind each class (Rationale Generation ).
+To ensure consistency, our TESOL-certificated professional linguist has developed an initial guideline inspired by Chen et al.
+(2020), which was also adopted by various well-known works Shon et al.
+(2022, 2023), and revised it frequently if necessary.
+Details of data annotation pipeline, annotation guidelines, data imbalance, translation annotation, and translaton quality control are shown in Appendix Section B.
+2.3 Data Quality Control
+During the independent annotation process conducted by three annotators, we observed a low inter-annotator agreement (Cohen’s kappa coefficient below 0.5 for the inter-annotator agreement between the two annotators), a common occurrence in real-world datasets as noted by Chen et al.
+(2020).
+To address this issue, we implemented an alternative label merging approach.
+We convened a discussion meeting involving the three annotators and two reviewers (one professional linguist and one with a biomedical background).
+Each annotator was required to justify their chosen sentiment label and its corresponding rationale.
+A label and its rationale were selected based on the consensus of all three annotators and two reviewers, rather than a majority vote, as employed in other studies Aziz and Dimililer (2020); Saleena et al.
+(2018).
 2.4 Data Statistics
@@
 Positive 524 20.01%
-Table 1: Distribution of sentiment labels in the dataset.
-Table 1 shows the distribution of sentiment labels in the dataset.
-In total, the dataset comprises 7878 samples with sentiment labels with rationales.
-This distribution reflects the dataset’s slight emphasis on neutral content, typical in medical conversations involving detailed explanations and advice.
+Table 1: Distribution of sentiment labels in the dataset for a single language.
+The real size of the dataset is 6 times larger when accounting all 5 languages - English, Chinese (Simplified and Traditional), German, and French.
+Dataset Venue #Samp. #Lang. Domain
+Mosi Zadeh et al. (2016) IEEE 3k 1 Vlog
+CMU-MOSEI Bagher Zadeh et al. (2018) ACL 23k 1 Various
+MELD Poria et al. (2019) ACL 13k 1 TV Series
+IEMOCAP Busso et al. (2008) Springer 12k 1 General
+SEMAINE McKeown et al. (2012) IEEE 1k 1 Simulation
+Sentiment Reasoning (ours) - 30k 5 Medical
+Table 2: Data statistics comparison based on the number of samples and languages.
+Our dataset with 5 languages (Vietnamese, English, Chinese, German and French) includes 30000 samples, making it the largest multimodal sentiment analysis dataset.
+Table 1 shows the distribution of sentiment labels in the dataset. This reflects the dataset’s slight emphasis on neutral content, typical in medical conversations involving explanations and advice.
+It should be noted that the statistics are reported for a single language, meaning that the real size of the dataset is 6 times larger when accounting all 5 languages.
 3 Sentiment Reasoning Framework
@@
 It comprises of two tasks: Sentiment Classification and Rationale Generation .
-Previous works Wadhwa et al.
-(2024); Chen et al.
-(2024); Hsieh et al.
-(2023); Ho et al.
-(2022) have shown that rationale-augmented targets consistently improve the performance of Language Models (LMs).
-Furthermore, having the option to output the model’s rationale during inference is highly desirable in many scenarios.
 3.2 Formal Definition
 Let $x^{T}_{1}:=x_{1},x_{2},...,x_{T}$ be an audio signal of length $T$.
-Let $C$ be the set of all possible sentiment classes, we should build a speech Sentiment Reasoning model $f$ that both estimates the probability $p(c|x^{T}_{1})$ for each $c\in C$ and generates its rationale sequence $r^{M}_{1}$ of $M$ length.
-The decision rule to predict a single sentiment class is:
+Let $C$ be the set of all possible sentiment classes, we should build a speech-based Sentiment Reasoning  model $f$ that both estimates the probability $p(c|x^{T}_{1})$ for each $c\in C$ and generates its rationale sequence $r^{M}_{1}$ of $M$ length.
+The decision rule to predict a sentiment class is:
 $x^{T}_{1}\to\hat{c}=\arg\max_{c\in C}f(c|x^{T}_{1})$ (1)
@@
 $w^{*}=\operatorname{arg}\max_{w_{1}^{N}}\,p(w_{1}^{N}|x_{1}^{T})$ (3)
-By utilizing Bayes’ Theorem, the probability $p(x)$ can be ignored during maximization since it merely serves as a normalization factor and does not affect the outcome:
-$\begin{split}p(w_{1}^{N}|x_{1}^{T})&=\frac{p(x_{1}^{T}|w_{1}^{N})p(w_{1}^{N})}{p(x_{1}^{T})}\\ &\propto p(x_{1}^{T}|w_{1}^{N})p(w_{1}^{N})\end{split}$ (4)
-Therefore:
-$w^{*}=\operatorname{arg}\max_{w_{1}^{N}}\underbrace{p(x_{1}^{T}|w_{1}^{N})}_{\text{acoustic model}}\cdot\underbrace{p(w_{1}^{N})}_{\text{language model}}$ (5)
 3.4 Language Model for Sentiment Reasoning
@@
 We employed hybrid ASR setup using wav2vec 2.0 encoder Le-Duc (2024) to transcribe speech to text.
-First, we generated alignments obtained by using Gaussian-Mixture-Model/Hidden-Markov-Model (GMM/HMM) as labels for wav2vec 2.0 Baevski et al.
-(2020) neural network training.
-The labels used in the acoustic modeling are context-dependent phonemes, triphones in this case.
-In GMM/HMM process, we used a CART (Classification And Regression Tree) Breiman (2017) to tie the states $s$, resulting 4501 CART labels:
-$\begin{split}&p(x_{1}^{T}|w_{1}^{N})=\sum_{[s_{1}^{T}]}\prod_{t=1}^{T}p(x_{t},s_{t}|s_{t-1},w_{1}^{N})\\ &=\sum_{[s_{1}^{T}]}\prod_{t=1}^{T}\underbrace{p(s_{t}|s_{t-1},w_{1}^{N})}_{\text{transition prob.}}\cdot\underbrace{p(x_{t}|s_{t},s_{t-1},w_{1}^{N})}_{\text{emission prob.}}\end{split}$ (8)
-After inputting CART labels for hybrid wav2vec 2.0 training, we employed frame-wise cross-entropy (fCE) loss Good (1952) to train the acoustic model.
-To transcribe speech given the acoustic observations, the acoustic model and n-gram language model Ney et al.
-(1994) should be combined based on the Bayes decision rule using Viterbi algorithm Forney (1973) which recursively computes the maximum path to a find best-path in the alignment graph of all possible predicted words to the acoustic observations:
-$\begin{split}w_{1}^{N}&=\operatorname{arg}\max_{N,w_{1}^{N}}p\Bigl{(}\prod_{n=1}^{N}p(w_{n}|w_{n-m}^{n-1})\\ &\cdot\max_{[s_{1}^{T}]}\prod_{t=1}^{T}p(x_{t},s_{t}|s_{t-1},w_{1}^{N})\Bigr{)}\end{split}$ (9)
-Finally, acoustic model and n-gram language model pruning (beam search) is used to only focus on the most promising predicted words at each time step $t$ Ortmanns et al. (1997).
-The final ASR model has 118M trainable parameters and Word-Error-Rate (WER) of 29.6% on VietMed test set.
+The final ASR model has 118M trainable parameters and Word-Error-Rate (WER) of 29.6% on the test set.
+Details of ASR experiments are shown in Appendix C.1.
+4.2 End-to-end Sentiment Classification
+We fine-tuned two well-known models, PhoWhisper Le et al.
+(2024) and Qwen2-Audio Chu et al.
+(2024), for the end-to-end spoken sentiment analysis task.
+PhoWhisper is trained large-scale ASR training set consisting of 844 hours of Vietnamese audio, while Qwen2-Audio is trained on more than 500 hours of audio.
+We use the base version of PhoWhisper with 74M parameters, while Qwen2-Audio has 8.2B parameters.
 4.3 Language Model for Sentiment Reasoning
@@
 However, encoders can not generate rationales.
-We use phoBERT (110M params) Nguyen and Nguyen (2020), a version of RoBERTa Liu et al.
+As such, they serve as baselines in our experiments.
+We use phoBERT (110M params) Nguyen and Nguyen (2020), RoBERTa Liu et al.
 (2019) pre-trained on 20GB Vietnamese text, and ViHealthBERT (110M params) Minh et al.
@@
 (2023a) as their backbone.
-These models were chosen based on their performance on the vmlu benchmark (Vietnamese Multitask Language Understanding) [^3] https://vmlu.ai/leaderboard which compares the performance of Vietnamese Large Language Models (LLMs) on various language understanding tasks.
+These models were chosen based on their performance on the vmlu benchmark (Vietnamese Multitask Language Understanding) [^2] https://vmlu.ai/leaderboard.
 4.4 Training with Rationale
+Previous works Wadhwa et al.
+(2024); Chen et al.
+(2024); Hsieh et al.
+(2023); Ho et al.
+(2022) have shown that rationale-augmented targets consistently improve the performance of generative language models.
 Our rationale-augmented training methods are based on, to our knowledge, the current state-of-the-art CoT-distillation approaches for each architecture.
@@
 (2024); Wadhwa et al.
-(2024) and compared to pre-thinking where the model first generates its chain-of-thought then provide the label (¡RATIONALE¿ ¡LABEL¿), post-thinking is more stable and token-efficient Chen et al.
+(2024) and compared to pre-thinking where the model first generates its CoT then provide the label (¡RATIONALE¿ ¡LABEL¿), post-thinking is more stable and token-efficient Chen et al.
 (2024); Wadhwa et al.
@@
 4.5 Rationale Format
-We further study the effects of the format of the rationale on the performance of the generative models.
-In particular, given the human rationale and human label, we further prompt GPT-3.5-turbo [^4] https://platform.openai.com/docs/models to enhance the rationale into two different format:
+While the rationale in our dataset were re-labeled by humans, we are also interested in whether a different and more detailed rationale format would help the models learn better.
+To this end, we further study the effects of the format of the rationale on the performance of the generative models.
+In particular, given the human rationale and human label, we further prompt GPT-3.5-turbo to enhance the rationale into two different format:
 Elaborated rationale: An elaborated version of the human rationale that is 1-2 sentence(s) long, grounded on the provided human rationale and the sentiment label.
@@
 4.6 Evaluation Metrics
-For Sentiment Classification task, we employ both accuracy and F1 score, the harmonic mean of precision and recall.
-It is particularly useful for evaluating models, as it considers both false positives and false negatives, providing a more comprehensive performance metric than accuracy alone.
-For Rationale Generation, we employ ROUGE (Recall-Oriented Understudy for Gisting Evaluation) score Lin (2004), which is a set of metrics used to evaluate the quality of generated text by comparing the overlap of n-grams Ney et al.
-(1994), word sequences, and word pairs between the generated texts and reference texts.
+For Sentiment Classification task, we employ accuracy and class-wise F1 score.
+For Rationale Generation , we employ ROUGE (Recall-Oriented Understudy for Gisting Evaluation) score Lin (2004).
 Also, we employ BERTScore Zhang et al.
-, which is an evaluation metric that leverages pre-trained BERT embeddings Devlin et al.
-(2019) to compute cosine similarity scores between generated texts and reference texts, capturing contextual and semantic nuances more effectively than traditional count-based metrics like ROUGE.
-BERTscore correlates well with human judgment on sentence-level as proved by Zhang et al.
-.
+which captures the contextual and semantic nuances.
+BERTscore has shown to correlate well with human judgment.
 5 Results and Analysis
@@
 Encoders are efficient yet effective Sentiment Classification baselines: Encoder models yields the best performance compared to their encoder-decoder and decoder counterparts, with high accuracy scores (¿ 0.665) and stable F1 scores (macro F1 of both models ¿ 0.665).
-Furthermore, compared to other models, encoders are more parameter efficient which makes them highly economical for training and deployment.
 We further observe that domain-specific encoders yield notably better performance, with ViHealthBERT outperforming phoBERT in accuracy (+0.8%) and macro F1 (+0.9%).
 2.
+ASR errors have a marginally negative impact on Sentiment Classification performance: For a fair comparison in real-world environments, WERs for human annotators on a standard conversational spontaneous English ASR dataset range from 5% to 15% Stolcke and Droppo (2017) while more challenging real-world ASR datasets are between 17% and 31% Mulholland et al.
+(2016).
+Given the complexity of real-world medical conversations, WER of 29.6% by our ASR model is within an acceptable range.
+Despite the WER of 29.6%, the performance drop in macro F1 scores is small (absolute value of only about 5%).
+3.
 Rationale-augmented training improve model performance: Consistent with previous findings, performing CoT-augmented training on both encoder-decoders and decoders improve our models performance compared to the baseline.
-Aside from ViT5, other models seem to have benefited from rationale-augmented training, especially the decoder-based LLMs.
-We observe an average accuracy gain of approximately +0.8% and average macro-F1 gain of approximately +1.3% for these models compared to their label-only counterparts.
 We further conducted a Student’s t-test Student (1908) and found that the gains are statistically significant for $\alpha=0.1$.
@@
 We observe a decline in all of our models performance on ASR data which is anticipated due to its WER of 29.6 %.
-Nonetheless, the models trained with rationale perform noticeably better than models without, with an average accuracy gain of +0.85% and macro F1 gain of +1.4%.
-3.
+Nonetheless, the models trained with rationale perform noticeably better than models without, with an average absolute accuracy gain of +0.85%, absolute macro F1 gain of +1.4%, and relative macro F1 gain of +2.5%.
+4.
 The format of post-thinking rationale doesn’t affect the generative models performance: We study the effects of the format of post-thinking rationale on the performance of generative models on Table 5 and observe that it is unclear whether there is a performance gain from more elaborated rationales.
@@
 (2024).
-Particularly, we observe that some models trained on the elaborated rationales have better performance (ViT5 and vmlu-llm).
-However, for the other models, there is a notable decline in all metrics.
-4.
+5.
 Models are likely to misclassify POSITIVE and NEGATIVE transcripts as NEUTRAL: We study the confusion matrix of our best model on human transcript, Vistral7B finetuned with human rationale, on Figure 2.
@@
 On the other hand, we found that models can easily distinguish NEGATIVE transcripts from POSITIVE ones.
-This reflects the ambiguity of sentiment analysis data, as in certain cases there is not enough distinctive features that differentiate neutral from other sentiments.
+This reflects the ambiguity of sentiment analysis data.
 Furthermore, given the slightly imbalanced nature of our dataset with fewer POSITIVE examples, its average F1 score is the lowest among the three labels across all models.
-5.
+6.
 Analysis of Generated Rationale: Compared to human rationale, we observe from Table 3 and Table 4 that the models trained with rationale have high BERTscore (around 0.8) with low ROUGE score, indicating that while the vocabulary used in the rationale is different, the overall semantic of the generated rationale remains similar to that of humans.
 Also, no noticeable changes in the semantic quality of rationale between human transcripts and ASR transcripts because BERTScore is still about 0.8 on both settings.
+Model Acc. F1 Neg. F1 Neu. F1 Pos. Mac F1
+PhoWhisper 0.4651 0.4393 0.5277 0.3328 0.4333
+Decoder (Label only)
+Qwen2-Audio 0.5815 0.5707 0.6150 0.5208 0.5688
+Decoder (Label + Rationale)
+Qwen2-Audio 0.5884 0.5875 0.6131 0.5337 0.5781
+Table 6: Performance of audio language models
+7.
+Results on end-to-end audio language models We report the results for end-to-end spoken sentiment analysis on PhoWhisper Le et al.
+(2024) and Qwen2-Audio Chu et al.
+(2024).
+Based on the results in Table 6, we make two observations: First, the performance of PhoWhisper is sub-optimal which we attribute to the fact that it was pre-trained for ASR-based tasks.
+Second, we found that rationale-augmented training can also increase the Sentiment Classification performance for audio language models.
 6 Conclusion
-In this work, we introduce a new task - Sentiment Reasoning - for both speech and text modalities, along with the framework and dataset.
-In Sentiment Reasoning, given an input transcript (either human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label).
+In this work, we introduce a new task - Sentiment Reasoning  - for both speech and text modalities, along with the framework and the world’s largest multimodal sentiment analysis dataset.
+In Sentiment Reasoning, given an input transcript (human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label).
 It comprises of two tasks: Sentiment Classification and Rationale Generation.
-We thoroughly evaluate the use of rationale during training to improve our models’ interpretability and performance.
-We found that rationale-augmented training improves model performance in Sentiment Classification in both human and ASR transcripts (1% increase in both accuracy and macro-F1).
-We also found that the generated rationales generally have different vocabulary to that of human but with similar semantics.
-Finally, we found no significant difference in the semantic quality of generated rationales between human and ASR transcripts.
+We menticulously evaluate the use of rationale during training to improve our models’ interpretability and performance.
+We found that rationale-augmented training improves model performance in Sentiment Classification in both human and ASR transcripts ( +2% increase in both accuracy and macro-F1).
+We found that the generated rationales have different vocabulary to human rationale but with similar semantics.
+Finally, we found no major difference in the semantic quality of generated rationales between human and ASR transcripts.
+7 Acknowledgement
+We thank Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for insightful feedback.
 8 Limitations
-This study utilized the hybrid ASR system, which is generally recognized as superior in performance compared to the attention-based encoder-decoder or end-to-end ASR systems Lüscher et al.
+Hybrid ASR: This study utilized the hybrid ASR system, which is generally recognized as superior in performance compared to the attention-based encoder-decoder or end-to-end ASR systems Lüscher et al.
 (2019); Prabhavalkar et al.
@@
 However, the hybrid ASR requires multiple steps, beginning with acoustic feature extraction and progressing through GMM-HMM modeling before transitioning to DNN-HMM modeling, which complicates reproducibility for non-experts.
-Additionally, we employed a cascaded speech Sentiment Reasoning approach.
-This approach uses a previously trained ASR model to generate transcripts that are subsequently input into a LM for downstream sentiment classification and reasoning tasks.
+Cascaded speech sentiment analysis approach: While we do report the results for end-to-end systems, our main focus in this paper is on cascaded speech sentiment analysis for Sentiment Reasoning  .
+This approach uses a previously trained ASR model to generate ASR transcripts that are subsequently input into a language model (LM) for downstream Sentiment Classification and Rationale Generation tasks.
 Consequently, the weights in the ASR model remain unchanged while the LM weights are updated.
-In spoken language processing, cascaded approach is prefered due to its straightforwardness, simplicity and superior accuracy Lu (2023); Bentivogli et al.
+In this setting, only semantic features from speech are utilized, omitting other trainable acoustic features, like prosody, tones, etc.
+In spoken language processing, where semantic features play a more important role than other acoustic features, cascaded approach is prefered due to its straightforwardness, simplicity and superior accuracy Lu (2023); Bentivogli et al.
 (2021); Tran et al.
@@
 (2023).
-Future work should consider the end-to-end sentiment analysis task, where weights in both the ASR model and LM are updated simultaneously, as it holds promise for improved performance.
-8 Acknowledgement
-We would like to thank Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for giving insightful feedback.
+Future work should consider the end-to-end sentiment analysis task, where weights in both the ASR model and LM are updated simultaneously, as it might hold promise for improved performance.
 Appendix A Related Works
@@
 (2020); Shon et al.
-(2021); Wu et al.
+(2021a); Wu et al.
 (2022); Tashev and Emmanouilidou (2019); Kaushik et al.
@@
 Appendix B Details about Data
-B.1 Data Annotation
+B.1 Data Annotation Pipeline
 We use LLM pre-labeling as it helps speed up the labeling process through providing the annotators with the initial sentiment labels and the corresponding rationales.
@@
 Existing works Chen et al.
-(2020) use both emotions and facts for sentiment labeling.
+(2020); Mohammad (2016); Shon et al.
+(2021b, 2022, 2023) use both emotions and facts for sentiment labeling.
+  • Emotion: Existing literature includes “emotion” as part of “sentiment” Chen et al.
+(2020); Shon et al.
+(2021b); Mohammad (2016) and sentiment analysis can be considered a more abstract level of emotion recognition, e.g.
+polarity of emotions Mohammad (2016).
+  • Facts: Many sentiment analysis systems require statements that describe events/situations to be given a sentiment label Chen et al. (2020); Mohammad (2016).
+The annotation task consists of two primary steps.
+First, annotators are required to perform Sentiment Classification.
+Second, annotators are instructed to provide a rationale behind each class (Rationale Generation ).
 To ensure consistency, our TESOL-certificated professional linguist has developed an initial guideline inspired by Chen et al.
-(2020) and revised it frequently if necessary as followed:
+(2020), which was also adopted by various well-known works Shon et al.
+(2022, 2023), and revised it frequently if necessary as followed:
+B.3.1 Output Annotation
 The NEGATIVE label is for chunks that discuss negative, serious diseases, disorders, symptoms, risks, negative emotions, or counter-positive statements (e.g.
@@
 Additionally, the presence of modal verbs (e.g., should, would, need) often indicates advice sharing, thereby classifying the chunk as NEUTRAL regardless of its content.
-B.4 Data Quality Control
-During the independent annotation process conducted by three annotators, we observed a low inter-annotator agreement, a common occurrence in real-world datasets as noted by Chen et al.
-(2020).
-To address this issue, we implemented an alternative label merging approach.
-We convened a discussion meeting involving the three annotators and two reviewers (one professional linguist and one with a biomedical background).
-Each annotator was required to justify their chosen sentiment label and its corresponding rationale.
-A label and its rationale were selected based on the consensus of all three annotators and two reviewers, rather than a majority vote, as employed in other studies Aziz and Dimililer (2020); Saleena et al.
-(2018).
-B.5 Data Statistics Discussion
-The NEUTRAL category is the most predominant, accounting for a significant portion of the dataset.
+B.4 Annotation Flowchart
+Inspired by the well-known annotation flowchart provided by Chen et al. (2020), we asked annotators to adopt the annotation flowchart and we ,if necessary, revised as follows:
+  1. 1.
+    Does the segment exhibit distinct emotional cues indicative of sentiment, such as laughter for positive affect or yelling for negative affect?
+    • Yes – Annotate the corresponding class and also note that:
+      • (a) In some instances, individuals may laugh to mitigate the discomfort associated with delivering negative statements. In such cases, it should be classified as neutral.
+      • (b) If individuals exhibit a sneer (a smile or laughter with a mocking tone), the corresponding sentiment should be classified as negative in such instances.
+    • No - Jump into Step 2
+  2. 2.
+    Does the segment provide an objective account of the facts?
+    • Yes - If the segment lists several positive attributes (e.g., good progress in medical treatment, good signs of health improvement), it is classified as positive.
+Conversely, if it lists several negative attributes, it is classified as negative.
+In the absence of a clear preponderance of either, the segment is considered neutral.
+    • No - Jump into Step 3
+  3. 3.
+    Does the segment exhibit a preference?
+    • Yes - If the subjective opinion or preference conveys a like or dislike, or expresses a positive (e.g., ”it is beneficial that…”) or negative sentiment, it should be annotated accordingly.
+    • No - It’s neutral
+  4. 4.
+    If the utterance is insufficient in length to accurately assess sentiment, it should be classified as neutral.
+B.5 Data Imbalance Discussion
+As shown in Table 1, NEUTRAL category is the most predominant, accounting for a significant portion of the dataset.
 With 3802 instances for both train and test set, NEUTRAL sentiments make up approximately half of the dataset.
@@
 POSITIVE sentiments typically involve discussions about recovery processes, positive outcomes, and favorable emotions.
-B.6 Data Samples
+A slight bias in the distribution of the labels towards NEUTRAL in our dataset (49.94% in the train set, 43.88% in the test set) reflects the nature of real-world medical conversations, rather than a weakness of our work.
+For context, in comparable real-world sentiment analysis datasets such as Switchboard-Sentiment Chen et al.
+(2020), the distribution is as follows: 30.4% of the speech segments are labelled as POSITIVE, 17% of the segments are labelled as NEGATIVE, and 52.6% of the segments are labelled as NEUTRAL.
+To address this labeling bias issue, future works can leverage techniques for fine-tuning models in data imbalance regimes, such as focal loss Ross and Dollár (2017), class weighting King and Zeng (2001).
+B.6 Translation Annotation Process and Translation Quality Control
+The data were initially translated from the source language into target languages (many-to-many) using the Gemini Large Language Model (LLM).
+Following the annotation process by zheng2023judging, the LLM-generated translated transcripts were treated as outputs from a real human annotator.
+In the data quality process, five human annotators manually corrected and then cross-verified all these translations based on the context of the whole conversation.
+Only transcripts that received consensus approval from multiple annotators were retained, resulting in an inter-annotator agreement of 100%.
+All human annotators possessed a professional language proficiency of C1 or higher (or HSK5 for Chinese) in their respective working languages.
+Additionally, each annotator had completed basic medical training and demonstrated substantial knowledge of medical terminology in their selected language.
+Furthermore, they were either currently pursuing or had completed undergraduate or graduate studies in countries where their chosen language is predominantly spoken.
+B.7 Data Samples
 Table 7 shows 9 examples with 3 samples per sentiment label in our dataset.
+As the Vietnamese transcripts are obtained from short-formed audio, the transcripts contain characteristics of spoken language which serve as noises to the model (e.g.
+stuttering, hesitation, etc).
+In our English translation, we aim to retain these properties, leading to unnatural, incomplete sentence with broken wording.
+Figure 3 shows 3 examples per sentiment label for all languages: Vietnamese, English, Chinese (Simplified and Traditional), German and French.
 Transcript ENG Translation Label Rationale
@@
 Table 7: 9 examples with 3 samples per sentiment label and its corresponding rationale
+Figure 3: Some samples from our dataset with versions all available languages.
 Appendix C Details about Experimental Setups
-C.1 Training Setup
+C.1 Details of ASR Experiments
+We employed hybrid ASR setup using wav2vec 2.0 encoder Le-Duc (2024) to transcribe speech to text.
+First, we generated alignments obtained by using Gaussian-Mixture-Model/Hidden-Markov-Model (GMM/HMM) as labels for wav2vec 2.0 Baevski et al.
+(2020) neural network training.
+The labels used in the acoustic modeling are context-dependent phonemes, triphones in this case.
+In GMM/HMM process, we used a CART (Classification And Regression Tree) Breiman (2017) to tie the states $s$, resulting 4501 CART labels:
+$\begin{split}&p(x_{1}^{T}|w_{1}^{N})=\sum_{[s_{1}^{T}]}\prod_{t=1}^{T}p(x_{t},s_{t}|s_{t-1},w_{1}^{N})\\ &=\sum_{[s_{1}^{T}]}\prod_{t=1}^{T}\underbrace{p(s_{t}|s_{t-1},w_{1}^{N})}_{\text{transition prob.}}\cdot\underbrace{p(x_{t}|s_{t},s_{t-1},w_{1}^{N})}_{\text{emission prob.}}\end{split}$ (6)
+After inputting CART labels for hybrid wav2vec 2.0 training, we employed frame-wise cross-entropy (fCE) loss Good (1952) to train the acoustic model.
+To transcribe speech given the acoustic observations, the acoustic model and n-gram language model Ney et al.
+(1994) should be combined based on the Bayes decision rule using Viterbi algorithm Forney (1973) which recursively computes the maximum path to a find best-path in the alignment graph of all possible predicted words to the acoustic observations:
+$\begin{split}w_{1}^{N}&=\operatorname{arg}\max_{N,w_{1}^{N}}p\Bigl{(}\prod_{n=1}^{N}p(w_{n}|w_{n-m}^{n-1})\\ &\cdot\max_{[s_{1}^{T}]}\prod_{t=1}^{T}p(x_{t},s_{t}|s_{t-1},w_{1}^{N})\Bigr{)}\end{split}$ (7)
+Finally, acoustic model and n-gram language model pruning (beam search) is used to only focus on the most promising predicted words at each time step $t$ Ortmanns et al. (1997).
+The final ASR model has 118M trainable parameters and Word-Error-Rate (WER) of 29.6% on VietMed test set.
+C.2 Training Setup
 Our encoders and encoder-decoders were trained on a cluster of 2 NVIDIA A40s with 46 GBs of memory.
@@
 Note that we do not perform hyperparameter tuning as we only aim to provide the initial benchmark results as well as studying the effects of CoT-augmented finetuning.
-C.2 Student’s T-Test
+C.3 Student’s T-Test
 A Student’s t-test, is a statistical method used to compare the means of one or two populations through hypothesis testing.
@@
 Appendix D Results on English subset
-We use Google Translate to translate our data into English.
-To ensure that our translated data is usable, we randomly sampled 50 transcripts and check their quality.
-We further train English models on this English subset of our dataset to ensure full usability.
+We randomly sampled 50 transcripts and check their quality. We further train English models on this English subset of our dataset to ensure full usability.
 The result of our experiments is in Table 8.
@@
 This finding is consistent with what when observed in our experiments in Section 5.
-Model Acc. F1 Neg. F1 Neut. F1 Pos. Mac F1
+Model Acc. F1 Neg. F1 Neu. F1 Pos. Mac F1
 Encoder (Label Only)
@@
 Flan-T5 Chung et al. (2022) 0.6326 0.6487 0.6390 0.5978 0.6285
-LLM (Label only)
+Decoder (Label only)
 Mistral7B Jiang et al. (2023b) 0.6290 0.6536 0.6322 0.5850 0.6236
-LLM (Label + Rationale)
+Decoder (Label + Rationale)
 Mistral7B Jiang et al. (2023b) 0.6454 0.6768 0.6364 0.6176 0.6436
 Table 8: Baseline performance of encoders, encoder-decoders, LLMs on the English human transcript. Further information about our metrics can be found in Table 2.
-Appendix E Error Analysis
+Appendix E Results on end-to-end audio language models
+E.1 Encoder-Based
+PhoWhisper Le et al.
+(2024): Whisper Radford et al.
+(2022) trained on an 844-hour Vietnamese dataset.
+It achieves the state-of-the-art performance on various Vietnamese ASR datasets.
+To perform Sentiment Classification, we attach a classification head to the encoder layer of the model.
+We use the PhoWhisper base version in our experiments.
+E.2 Audio LLMs
+Qwen2-Audio Chu et al.
+(2024): the state-of-the-art Qwen large audio-language model with 7B parameters, pretrained on diverse audio datasets.
+We use the Instruct version of this model.
+We finetune Qwen2-Audio on both Label Only and Label + Rationale settings in our experiments.
+It should be note that the rationales are text-only.
+Appendix F Error Analysis
 Transcript Label Pred. Human Rationale Model Rationale

### v5

@@
 (2022) have revealed that training generative small language models (SLMs) on rationale-augmented targets (the CoT from larger models is provided along side with the target label) can help the SLM (1) perform better and (2) acquire the ability to generate rationale.
-Our work leverage these findings and prepare a set of human-labeled rationale to train our sentiment analysis models to do Rationale Generation and enhance their performance (Section 4.4 and 4.5).
+Our work leverage these findings and prepare a set of human-labeled rationale to train our sentiment analysis models to do Rationale Generation and enhance their performance (Section 4.5 and 4.6).
 By incorporating rationales to explain reason for each predicted sentiment label, users could understand the model’s reasoning, facilitating better decision-making based on the classification results.
@@
 $w^{N}_{1}\to\hat{c}=\arg\max_{c\in C}g(c|w^{N}_{1})$ (4)
-$g$ is trained to minimize a loss function $\mathscr{L}(g(w^{N}_{1}),\hat{c})$.
-The optimal parameters $\theta$ of the model are found by solving the optimization problem $\min_{\theta}\mathscr{L}(g(w^{N}_{1};\theta),\hat{c})$.
-Once trained, the model can predict the class of the transcribed audio signal by evaluating $\hat{c}=g(w^{N}_{1})$.
+[…]
 3.4.2 Rationale Generation
@@
 4.2 End-to-end Sentiment Classification
-We fine-tuned two well-known models, PhoWhisper Le et al.
-(2024) and Qwen2-Audio Chu et al.
-(2024), for the end-to-end spoken sentiment analysis task.
-PhoWhisper is trained large-scale ASR training set consisting of 844 hours of Vietnamese audio, while Qwen2-Audio is trained on more than 500 hours of audio.
-We use the base version of PhoWhisper with 74M parameters, while Qwen2-Audio has 8.2B parameters.
+PhoWhisper Le et al.
+(2024): The encoder-based Whisper Radford et al.
+(2022) trained on an 844-hour Vietnamese dataset.
+To perform Sentiment Classification, we attach a classification head to the encoder layer of the model.
+We use the base version in our experiments.
+4.3 End-to-end Sentiment Reasoning
+Qwen2-Audio Chu et al. (2024): We fine-tune the Qwen2-Audio (7B parameters) on both Label Only and Label + Rationale settings. We use the Instruct version in our experiments.
 4.4 Language Model for Sentiment Reasoning
@@
 4.4.2 Generative Models
-[…]
+We reformulated Sentiment Classification into a text-to-text problem, where given the input transcript $w^{N}_{1}$, the generative model $g$ and the predicted sentiment class $c$, we have $g(w^{N}_{1})=c$ with $c\in C=\{\textit{"0", "1", "2"}\}$ where 0,1,2"0","1","2" corresponds to the labels NEGATIVE, NEUTRAL and POSITIVE.
 Encoder-Decoder: BARTpho (139M params) Tran et al.
@@
 7 Acknowledgement
-We thank Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for insightful feedback.
+We thank Professor Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for their insightful feedback.
 8 Limitations
@@
 The data were initially translated from the source language into target languages (many-to-many) using the Gemini Large Language Model (LLM).
-Following the annotation process by zheng2023judging, the LLM-generated translated transcripts were treated as outputs from a real human annotator.
+Following the annotation process by Zheng et al.
+(2023), the LLM-generated translated transcripts were treated as outputs from a real human annotator.
 In the data quality process, five human annotators manually corrected and then cross-verified all these translations based on the context of the whole conversation.
@@
 Our encoders and encoder-decoders were trained on a cluster of 2 NVIDIA A40s with 46 GBs of memory.
-All models were trained on 30 epochs with with a learning rate of $-52e\text{-}$ 5 and batch size of 64.
+All models were trained on 30 epochs with with a learning rate of $2e\text{-}5$ and batch size of 64.
 We evaluated every epoch with early stopping with patience = 3.
@@
 Table 8: Baseline performance of encoders, encoder-decoders, LLMs on the English human transcript. Further information about our metrics can be found in Table 2.
-Appendix E Results on end-to-end audio language models
-E.1 Encoder-Based
-PhoWhisper Le et al.
-(2024): Whisper Radford et al.
-(2022) trained on an 844-hour Vietnamese dataset.
-It achieves the state-of-the-art performance on various Vietnamese ASR datasets.
-To perform Sentiment Classification, we attach a classification head to the encoder layer of the model.
-We use the PhoWhisper base version in our experiments.
-E.2 Audio LLMs
-Qwen2-Audio Chu et al.
-(2024): the state-of-the-art Qwen large audio-language model with 7B parameters, pretrained on diverse audio datasets.
-We use the Instruct version of this model.
-We finetune Qwen2-Audio on both Label Only and Label + Rationale settings in our experiments.
-It should be note that the rationales are text-only.
-Appendix F Error Analysis
+Appendix E Error Analysis
+We report our best model’s misclassified transcripts with the highest label confidence (defined as the softmax of the logits of the model prediction) in Table 9.
+By analyzing at the model’s rationale, we hypothesize that the model is confounded by the appearance of certain keywords that elicit either extremely positive ( hữu ích (helpful)) or negative, disease-related words and sentiment which pushes the model away from the NEUTRAL label.
 Transcript Label Pred. Human Rationale Model Rationale
@@
 Table 9: Some misclassified transcripts from our best model with high confidence (¿0.99). VI means the Vietnamese transcript, EN means the transcript translated to English
-We report our best model’s misclassified transcripts with the highest label confidence (defined as the softmax of the logits of the model prediction) in Table 9.
-By analyzing at the model’s rationale, we hypothesize that the model is confounded by the appearance of certain keywords that elicit either extremely positive ( hữu ích (helpful)) or negative, disease-related words and sentiment which pushes the model away from the NEUTRAL label.